# version 0.6

### v3

@@ 7.5 Exercises
Chapter 8 Policy Classes for Partially Observed LDS
In this chapter we consider linear time varying systems with partial observation, i.e. the state is not available to the controller. Instead, a (known or unknown) linear projection of the state is revealed. In full generality, recall from definition 1.1 partially observed dynamical system is given by the equations
As per our notation thus far, here $\mathbf{x}_{t}\in{\mathbb{R}}^{d_{x}}$ is the state of the system, $\mathbf{y}_{t}\in{\mathbb{R}}^{d_{y}}$ is the observation, $\mathbf{u}_{t}\in{\mathbb{R}}^{d_{u}}$ is the control input, and $\mathbf{w}_{t}\in{\mathbb{R}}^{d_{w}}$ is the perturbation. The function $g_{t}$ is a mapping from the state to the observed state, and the function $f_{t}$ is the transition function, given current state $\mathbf{x}_{t}$ and control input $\mathbf{u}_{t}$. The subscript indicates the relevant quantities at time step $t.$
Many of the control policies we consider in this chapter are general and can be applied to any dynamical system of the above formulation. However, following the paradigm of chapter 6, we relate the representation power of policy classes only for partially observed linear dynamical systems. These are given by the equations
$\mathbf{x}_{t+1}$ $=A_{t}\mathbf{x}_{t}+B_{t}\mathbf{u}_{t}+\mathbf{w}_{t}$
$\mathbf{y}_{t}$ $=C_{t}\mathbf{x}_{t}.$
Often, partially observed LDSs are defined with an additional noise term $\mathbf{v}_{t}$ corrupting the observations as $\mathbf{y}_{t}=C_{t}\mathbf{x}_{t}+\mathbf{v}_{t}$. Exercise 8.5 asks the reader to prove that even in such circumstances we may assume the previously stated form for partially observed systems without loss of any generality.
We limit the discussion in this chapter to stable partially observed systems, although the policy containment results described next hold more broadly.
8.1 Linear Observation Action Controllers
Definition 8.1.
A linear observation-action controller $K\in{\mathbb{R}}^{d_{u}\times d_{y}}$ is a linear operator mapping observation to control as $\mathbf{u}_{t}^{\pi}=K\mathbf{y}_{t}$. We denote by $\Pi_{\gamma}^{OAC}$ the set of all linear controllers bounded in Fronbenius norm by $\|K\|\leq\gamma$, that are also $\gamma$-stabilizing for given LDS.
While this class of policies is natural to consider, it turns out to be very inexpressive. Even for optimal control with stochastic noise, the optimal policy cannot be approximated by a linear observation-action policy (see Exercise 8.5). Therefore we proceed to define more expressive and useful policies. However, OACs will be useful in simplifying proofs later on.
8.1.1 Generalized Linear Observation Action Controllers
Similar to the full observation setting, we can generalize linear observation-action controllers to apply to a history of past observations. This additional generality, of taking a history of observations, greatly increases the expressive power of this policy class. The formal definition is below.
Definition 8.2.
Parameterized by matrices $K_{0:h}$, generalized linear observation-action controllers choose a control input as
$\mathbf{u}_{t}^{\pi}=\sum_{i=0}^{h}K_{i}\mathbf{y}_{t-i}.$
Let $\Pi^{GOAC}_{h,\gamma}$ be that class of all such policies satisfying $\sum_{i=0}^{h}\|K_{i}\|\leq\gamma$ that are also $\gamma$-stabilzing.
Generalized observation-action policies can be seen as a special case of observation-action controllers using the technique of lifting from Section 6.2.3. Consider any partially-observed system $(A_{t},B_{t},C_{t})$ and perturbation sequence $\mathbf{w}_{t}$. Consider the lifted system $(\tilde{A}_{t},\tilde{B}_{t},\tilde{C}_{t},\tilde{D}_{t})$ given by
$\mathbf{z}_{t+1}=\tilde{A}_{t}\mathbf{z}_{t}+\tilde{B}_{t}\mathbf{u}_{t}+\tilde{D}_{t}\mathbf{w}_{t},\quad\tilde{\mathbf{y}}_{t}=C_{t}\mathbf{z}_{t},$
where
$\tilde{A}_{t}=\begin{array}[]{|c|c|c|c|c|}\hline\cr A_{t}&0&\ldots&0&0\\ \hline\cr I&0&\ldots&0&0\\ \hline\cr 0&I&\ldots&0&0\\ \hline\cr\vdots&\vdots&\cdots&\vdots&\vdots\\ \hline\cr 0&0&\cdots&I&0\\ \hline\cr\end{array}\ \ ,\ \ \tilde{B}_{t}=\begin{array}[]{|c|}\hline\cr B_{t}\\ \hline\cr 0\\ \hline\cr 0\\ \hline\cr\vdots\\ \hline\cr 0\\ \hline\cr\end{array}\ \ ,$
$\ \ \tilde{C}_{t}=\begin{array}[]{|c|c|c|c|}\hline\cr C_{t}&C_{t-1}&\dots&C_{t-h}\\ \hline\cr\end{array}\ \ ,\ \ \tilde{D_{t}}=\begin{array}[]{|c|}\hline\cr I\\ \hline\cr 0\\ \hline\cr 0\\ \hline\cr\vdots\\ \hline\cr 0\\ \hline\cr\end{array}.$
The following lemma relates the original and lifted system (for a proof see exercise 8.5).
Lemma 8.3.
8.2 Linear Dynamic Controllers
The main motivation for the class of linear dynamical controllers, which we defined already in the context of full observability, is that of partially observed linear dynamical systems. The generalization of this class is as follows.
Note that LDCs are $\varepsilon$-approximated by generalized observation-action controllers, namely GOACs. The proof is left as Exercise 9.4, on account of bering similar to that for the fully observed case.
Lemma 8.5.
Consider a partially-observed time-invariant linear dynamical system $\{A,B,C\}$. The class $\Pi^{GOAC}_{h^{\prime},\gamma^{\prime}}$ $\varepsilon$-approximates the class $\Pi^{LDC}_{\gamma}$ for $h^{\prime}=O\left(\gamma\log\frac{\gamma}{\varepsilon}\right),\gamma^{\prime}=2\gamma^{3}$.
8.3 Disturbance Response Controllers
The most important policy class in the context of nonstochastic control is that of Disturbance Response Controllers, or DRC. These policies act on a signal called Nature’s y’s, denoted $\mathbf{y}^{\mathbf{nat}}_{t}$, which is the would-be observation at time $t$ assuming all inputs to the system were zero from the beginning of time, or formally,
$\mathbf{x}^{\mathbf{nat}}_{t+1}=A_{t}\mathbf{x}^{\mathbf{nat}}_{t}+\mathbf{w}_{t}\ ,\ \mathbf{y}^{\mathbf{nat}}_{t}=C_{t}\mathbf{x}_{t}.$
A linear disturbance response policy $\pi$ is parameterized by the matrices $M_{1:h}=[{M}_{0},\ldots,{M}_{h}]$, and outputs control $\mathbf{u}_{t}^{\pi}$ at time $t$ according to the rule
$\mathbf{u}_{t}^{\pi}=\sum_{i=0}^{h}{M}_{i}\mathbf{y}^{\mathbf{nat}}_{t-i}.$
Denote by $\Pi^{DRC}_{h,\gamma}$ the set of all disturbance response policies with $K_{1:T}$ being $\gamma$-stabilizing linear controllers for the system and
$\sum_{i=0}^{h}\|M_{i}\|\leq\gamma.$
8.3.1 The Expressive Power of Disturbance Response Controllers
A DRC always produces controls that ensure the system remains state-bounded. This can be shown analogously to Lemma 6.8 (see Exercise 8.5).
To demonstrate that DRCs approximately capture GOACs, we instead prove that OACs, a weaker class of policies, are $\varepsilon$-approximated by DRCs. Nevertheless, this observation is sufficient to conclude the stronger claim, given that GOACs can be written as OACs on a lifted system as noted in Lemma 8.3, and simultaneously that a DRC with history $H^{\prime}$ on the lifted system can be written as a DRC with history $H^{\prime}+h$ over the original system.
Lemma 8.6.
Consider a $\gamma$-stable time-invariant linear dynamical system $(A,B,C)$. The class $\Pi^{DRC}_{h^{\prime},\gamma^{\prime}}$ $\varepsilon$-approximates the class $\Pi^{OAC}_{\gamma}$ for parameters $h^{\prime}=O\left(\gamma\log\frac{\gamma}{\varepsilon}\right)$, $\gamma^{\prime}=O(\gamma^{6})$.
Proof.
Recall that for the fully observed case the state-action pairs produced by execution of a linear policy can be written as linear functions of the perturbations. It is this observation that allowed us to prove that DACs capture linear policies for fully observed systems. We prove an analogue for partially observed systems: The observation-action pairs produced by the execution of OACs can be written as linear functions of nature’s $y$’s. Thereafter, we argue that parameterizing the controllers to be linear in nature’s $y$’s (as DRCs do) allows us to encompass OACs.
We introduce an augmented partially observed linear system that helps us write the observation-action pairs produced by a OAC as a linear function of nature’s $y$’s. Crucially, the new system has the same nature’s $y$’s as the original system.
Lemma 8.7.
For any partially-observed system $(A_{t},B_{t},C_{t})$ and perturbation sequence $\mathbf{w}_{t}$ producing observations $\mathbf{y}_{t}$, construct a lifted system $(\tilde{A}_{t},\tilde{B}_{t},\tilde{C}_{t})$ given by
$\begin{bmatrix}\mathbf{x}^{\mathbf{nat}}_{t+1}\\ \Delta_{t+1}\end{bmatrix}=\begin{bmatrix}A_{t}&0\\ 0&A_{t}\end{bmatrix}\begin{bmatrix}\mathbf{x}^{\mathbf{nat}}_{t}\\ \Delta_{t}\end{bmatrix}+\begin{bmatrix}0\\ B_{t}\end{bmatrix}\mathbf{u}_{t}+\begin{bmatrix}I\\ 0\end{bmatrix}\mathbf{w}_{t},\quad\tilde{\mathbf{y}}_{t}=\begin{bmatrix}C_{t}&C_{t}\end{bmatrix}\begin{bmatrix}\mathbf{x}^{\mathbf{nat}}_{t}\\ \Delta_{t}\end{bmatrix}.$
Substituting $\mathbf{u}_{t}^{\pi}=K\mathbf{y}^{\pi}_{t}=KC(\mathbf{x}^{\mathbf{nat}}_{t}+\Delta^{\pi}_{t})$, the state-action sequence produced by a OLIN $\mathbf{u}_{t}^{\pi}=K\mathbf{y}_{t}^{\pi}$ can be written as
$\mathbf{u}^{\pi}_{t}$ $=K\mathbf{y}^{\mathbf{nat}}_{t}+KC\Delta^{\pi}_{t}$
$=\mathbf{u}_{t}+Z_{t}.$
Given $\gamma$-stabilizability of the OAC $K$ and $\gamma$-intrinsic stability of the of the system $(A,B,C)$, we have that for some $\kappa,\delta$ suc that $\frac{\kappa}{\delta}\leq\gamma$ that
The approximation in terms of observations and controls implies $\varepsilon$-approximation of policy class for any Lipchitz cost sequence, concluding the proof. ∎
8.4 Bibliographic Remarks
The Disturbance Response Parametrization was introduced in [SSH20] as an analogue of DAC for control of parially observed systems. It can be viewed as a succinct state-space representation of the Youla parameterization [YJB76]. For time-varying dynamics, the power of vaious policy classes was contrasted in [MGSH21].
8.5 Exercises
$\mathbf{x}_{t+1}$ $=A_{t}\mathbf{x}_{t}+B_{t}\mathbf{u}_{t}+\mathbf{w}_{t},$
$\mathbf{y}_{t}$ $=C_{t}\mathbf{x}_{t}+\mathbf{v}_{t}.$
$\mathbf{y}^{\prime}_{t}$ $=C^{\prime}_{t}\mathbf{z}_{t}$
8.2. Prove Lemma 8.3.
8.3. For any $\gamma$-stable partially observed system, prove that any DRC policy in $\Pi_{H,\gamma}^{DRC}$ is $\gamma^{2}$-stabilizing.
8.4. Construct a partially observed linear dynamical system subject to Gaussian i.i.d. perturbations with quadratic costs such that simultaneously it is true that (a) there exists a GOAC policy that incurs zero (or at most constant) aggregate cost, (b) yet any OAC policy incurs $\Omega(T)$ aggregate cost when run for $T$ time steps.
Chapter 9 Online Nonstochastic Control with Partial Observation
In this chapter we explore the more general setting of nonstochastic control that allows for partial observation. Partial observation introduces significant challenges and requires revisiting the controllability and stabilizability definitions we have discussed in previous chapters. To allow for more streamlined and simple exposition we make a significant simplification: we assume that the underlying linear dynamical system is stable. The generalization of the results presented in this chapter to stabilizable systems is surveyed in the bibliographic section at the end of this chapter.
Hence, we consider time varying linear dynamical systems with partial observation governed by the equations  [^1] Recall that this is a slightly simplified form, but without loss of generality, as per Exercise 8.5.
$\mathbf{x}_{t+1}$ $=A_{t}\mathbf{x}_{t}+B_{t}\mathbf{u}_{t}+\mathbf{w}_{t}$
$\mathbf{y}_{t}$ $=C_{t}\mathbf{x}_{t}.$
The definition of policy regret 1.3 naturally extends to partial observation as below. The setting is similar: the controller iteratively chooses a control $\mathbf{u}_{t}$, and then observes the next observation of the system $\mathbf{y}_{t+1}$ and suffers a loss of $c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})$, according to an adversarially chosen loss function. The states $\mathbf{x}_{t}$ is hidden from the controller.
Analogous to policy regret definition 1.3 aa Let $\Pi=\{\pi:\mathbf{y}_{1:t}\mapsto\mathbf{u}_{t}\}$ be a class of policies. The regret of the controller w.r.t. $\Pi$ is defined as follows.
Definition 9.1.
Let $\{A_{t},B_{t},C_{t}\}$ be a time varying, partially observable, linear dynamical system. The regret of an online control algorithm over $T$ iterations w.r.t. class of policies $\Pi$ is given by:
$\mathrm{{regret}}_{T}({\mathcal{A}},\Pi)=\max_{\xi_{1:T},\mathbf{w}_{1:T}:\mathopen{\|}\mathbf{w}_{t}\mathclose{\|},\|\xi_{t}\|\leq 1}$ $\left(\sum_{t=1}^{T}c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})-\min_{\pi\in\Pi}\sum_{t=1}^{T}c_{t}(\mathbf{y}_{t}^{\pi},\mathbf{u}_{t}^{\pi}))\right),$
where $\mathbf{u}_{t}={\mathcal{A}}(\mathbf{y}_{t})$ are the controls generated by ${\mathcal{A}}$, and $\mathbf{y}_{t}^{\pi},\mathbf{u}_{t}^{\pi}$ are the counterfactual state sequence and controls under the policy $\pi$, i.e.
$\mathbf{u}_{t}^{\pi}$ $=\pi(\mathbf{y}_{1:t}^{\pi})$
$\mathbf{x}_{t+1}^{\pi}$ $=A_{t}\mathbf{x}_{t}^{\pi}+B_{t}\mathbf{u}_{t}^{\pi}+\mathbf{w}_{t}$
$\mathbf{y}_{t}^{\pi}$ $=C_{t}\mathbf{x}_{t}^{\pi}.$
Henceforth, if $T$ and $\Pi,{\mathcal{A}}$ are clear from the context, we just refer to $\mathrm{{regret}}$ without quantifiers.
9.1 Disturbance Response Controllers
Recall from section 8 the policy class of disturbance response controllers. This is the natural analogue of disturbance action controllers from the full information setting, adapted to use a different signal called “Nature’s $\mathbf{y}$’s”.
Recall the simplification that we make in this chapter: the linear dynamical system is stable. This allows us to ignore the stabilizing controller component of disturbance action control and greatly simplifies the exposition of the main ideas. We thus can define in succinct notation,
Definition 9.2 (Disturbance Response Controller).
A disturbance response policy for stable systems $\pi_{M_{1:h}}$ is parameterized by the matrices $M_{1:h}=[{M}_{1},\ldots,{M}_{h}]$. It outputs control $\mathbf{u}_{t}^{\pi}$ at time $t$ according to the rule
$\mathbf{u}_{t}^{\pi}=\sum_{i=1}^{h}{M}_{i}\mathbf{y}^{\mathbf{nat}}_{t-i},$
Here $\Phi_{t}$ is the linear operator such that $\Phi_{t}(t-i)=\prod_{j=0}^{i-1}A_{t-j}$.
The main advantage of this policy class, as opposed to those considered in optimal control such as linear dynamical controllers, is that they yield convex parametrization. Thus, this policy class allows for provably efficient optimization and algorithms.
We conclude with the following important lemma on an expression that is important to compute nature’s y’s. The proof of this lemma follows directly from the definition and is left as exercise 9.4.
Lemma 9.3.
$\mathbf{z}_{t+1}$ $=A_{t}\mathbf{z}_{t}+B_{t}\mathbf{u}_{t},$
$\mathbf{y}^{\mathbf{nat}}_{t}$ $=\mathbf{y}_{t}-C_{t}\mathbf{z}_{t}.$
9.2 The Gradient Response Controller
Partial observation requires modification to the Gradient Perturbation Controller we have studied in the previous chapter. The Gradient Response Controller algorithm is described in figure (4), for general changing linear dynamical systems with partial observation. Similar to the GPC, the main idea is to learn the parametrization of a disturbance response control policy using known online convex optimization algorithms, such as online gradient descent. We henceforth prove that this parametrization is essentially convex, allowing us to prove the main regret bound.
In algorithm (4) we denote $\mathbf{y}_{t}({M}_{1:h})$ as the observation arising by playing DRC policy ${M}_{1:h}$ from the beginning of time, and same for $\mathbf{u}_{t}({M}_{1:h})$. Although it may be impossible to determine $\mathbf{w}_{t}$, especially if $C_{t}$ has fewer rows than columns, from the observations $\mathbf{y}_{t}$, $\mathbf{y}_{t}({M}_{1:h})$ can be determined from $\mathbf{y}^{\mathbf{nat}}_{t}$ alone by the application of Lemma 9.3, as
1:  Input: sequence $\{A_{t},B_{t},C_{t}\}$ , $h$, $\eta$, initialization $M_{0:h}^{1}$.
2:  for $t$ = $1\ldots T$ do
4:     Observe $\mathbf{y}_{t+1}$, compute for $t+1$:
5:     Construct loss $\ell_{t}({M}_{0:h})=c_{t}(\mathbf{y}_{t}({M}_{0:h}),\mathbf{u}_{t}({M}_{0:h}))$Update $M_{0:h}^{t+1}\leftarrow{M}_{0:h}^{t}-\eta\nabla\ell_{t}({M}_{0:h}^{t})$
6:  end for
Algorithm 4 Gradient Response Controller(GRC)
Theorem 9.4.
Assuming that
  • a
  • b
    The matrices $\{{A}_{t},{B}_{t},C_{t}\}$ have bounded $\ell_{2}$ norm.
  • c
    The LDS $\{{A}_{t},{B}_{t},C_{t}\}$ is $\gamma$-stable.
Then the GRC algorithm (4) ensures that
$\max_{\xi_{1:T},\mathbf{w}_{1:T}:\mathopen{\|}\mathbf{w}_{t}\mathclose{\|},\|\xi_{t}\|\leq 1}\left(\sum_{t=1}^{T}c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})-\min_{\pi\in\Pi^{DRC}}\sum_{t=1}^{T}c_{t}(\mathbf{y}_{t}^{\pi},\mathbf{u}_{t}^{\pi})\right)\leq\mathcal{O}(\sqrt{T}).$
The proof is largely analogous to the fully observed case. We present a sketch below.
Proof sketch.
As per Theorem A.1 in Appendix A, the online gradient descent algorithm for convex loss functions ensures that the regret is $\mathcal{O}(\sqrt{T})$. However, to directly apply the OGD algorithm to our setting, we need to make sure that the following 3 conditions hold true:
  • The loss function should be a convex function in the variables ${M}_{0:h}$.
  • The loss function is time dependent because the loss depends on the current observation/state, which further depends on the previous states. The state at each time step depends on the current variable estimates. Hence, $\mathbf{y}_{t}$ can be shown to be a function of $\left\{{M}^{i}_{0:h}\right\}_{i=1}^{t}$, where ${M}^{i}_{0:h}$ denotes the variables at step $i$. Thus, we need to make sure that optimizing the loss function is similar to optimizing a function which doesn’t have such a dependence on time.
Lemma 9.5.
The loss functions $\ell_{t}$ are convex in the variables ${M}_{0:h}$.
Proof.
Using $\mathbf{x}_{t}$ to denote $\mathbf{x}_{t}(M_{0:h})$ (and similarly for $\mathbf{y}_{t},\mathbf{u}_{t}$), the loss function $\ell_{t}$ is given by
$\ell_{t}({M}_{0:h})=c_{t}(\mathbf{x}_{t},\mathbf{y}_{t}).$
Since the cost $c_{t}$ is a convex function w.r.t. its arguments, we simply need to show that $\mathbf{y}_{t}$ and $\mathbf{u}_{t}$ depend linearly on ${M}_{0:h}$. Using Lemma 9.3, the action-observation sequence is given by
$\mathbf{u}_{t}$ $=\sum_{i=0}^{h}M_{i}\mathbf{y}^{\mathbf{nat}}_{t-i},$
$\mathbf{y}_{t}$ $=\mathbf{y}^{\mathbf{nat}}_{t}+C_{t}\sum_{i=1}^{t}\left[\prod_{j=1}^{i-1}{A}_{t-j}\right]B_{t-i}\mathbf{u}_{t-i}$
$=\mathbf{y}^{\mathbf{nat}}_{t}+C_{t}\sum_{i=1}^{t}\left[\prod_{j=1}^{i-1}{A}_{t-j}\right]B_{t-i}\sum_{j=0}^{h}M_{j}\mathbf{y}^{\mathbf{nat}}_{t-i-j}$
Now, by the regret bound implied by OGD, we have
$\sum_{t=1}^{T}l_{t}(M_{0:h}^{t})-\min_{M_{0:h}}\sum_{t=1}^{T}l_{t}(M_{0:h})\leq\mathcal{O}(\sqrt{T}).$ (9.1)
Finally, the actual loss $c_{t}$ at time $t$ is not calculated on $\mathbf{y}_{t}(M_{0:h})$, but rather on the true observation $\mathbf{y}_{t}$, which in turn depends on different parameters $M_{0:h}^{i}$ for various historical times $i<t$. Nevertheless, $c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})$ is well approximated by $l_{t}(M^{t}_{0:h})$.
Lemma 9.6.
For all $t$, $|l_{t}(M^{t}_{0:h})-c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})|\leq\mathcal{O}(1/\sqrt{T})$.
Proof.
First, note that to achieve $\mathcal{O}(\sqrt{T})$ regret OGD uses a step size of $\eta=1/\sqrt{T}$. Since the gradient of the loss functions are bounded, we have
$\|M_{0:h}^{t}-M_{0:h}^{t-i}\|\leq\sum_{s=t-i+1}^{t}\|M_{0:h}^{s}-M_{0:h}^{s-1}\|\leq\mathcal{O}\left(\frac{i}{\sqrt{T}}\right).$
Now, we use the fact that the iterates of OGD $M_{1:h}^{t}$ move slowly to establish that $\mathbf{y}_{t}$ and $\mathbf{y}_{t}(M^{t}_{0:h})$ are close in value. Before that, by definition, we have $\mathbf{u}_{t}(M^{t}_{0:h})=\mathbf{u}_{t}$ since the sequence $\mathbf{y}^{\mathbf{nat}}_{t}$ is determined independently of the choice of the policy being executed. Due to the $\gamma$-stability of the underlying system, we have for any $i\leq j$, $\|C_{j}\prod_{t=i+1}^{j-1}{A}_{t}B_{i}\|\leq\kappa(1-\delta)^{j-i}$ for some $\kappa>0,\gamma<1$. Using Lemma 9.3, we have for some $C>0$ that
$\|\mathbf{y}_{t}(M^{t}_{0:h})-\mathbf{y}_{t}\|$ $=\left\|C_{t}\sum_{i=1}^{t}\left[\prod_{j=1}^{i-1}A_{t-j}\right]B_{t-i}(\mathbf{u}_{t-i}(M^{t}_{0:h})-\mathbf{u}_{t-i})\right\|$
$\leq\sum_{i=1}^{t}\left\|C_{t}\left[\prod_{j=1}^{i-1}A_{t-j}\right]B_{t-i}\right\|\|(\mathbf{u}_{t-i}(M^{t}_{0:h})-\mathbf{u}_{t-i})\|$
$\leq C\kappa\sum_{i=1}^{t}(1-\delta)^{i}\frac{i}{\sqrt{T}}\max_{j}\|\mathbf{y}^{\mathbf{nat}}_{j}\|$
$\leq\frac{C\kappa}{\delta^{2}\sqrt{T}}\max_{j}\|\mathbf{y}^{\mathbf{nat}}_{j}\|\leq\frac{C\kappa^{2}}{\delta^{3}\sqrt{T}}$
Noting that $c_{t}$ has bounded derivatives concludes the proof of the lemma. ∎
Now, we combine the regret guarantee in equation 9.1 and the immediately preceding lemma as
$\sum_{t=1}^{T}c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})-\min_{\pi\in\Pi^{DRC}}\sum_{t=1}^{T}c_{t}(\hat{\mathbf{y}}_{t},\pi(\hat{\mathbf{y}}_{0:t}))$
$\leq\sum_{t=1}^{T}l_{t}(M_{0:h}^{t})-\min_{\pi\in\Pi^{DRC}}l_{t}(M_{0:h}^{\pi})+T\times\mathcal{O}\left(\frac{1}{\sqrt{T}}\right)=\mathcal{O}(\sqrt{T}),$
implying the stated claim.
9.3 Bibliographic Remarks
9.4 Exercises
9.1. Prove Lemma 9.3.
9.2. Consider a partially observed linear dynamical system that is $\gamma$-stable in the absence of control inputs. Prove that when a DRC policy $\pi\in\Pi^{DRC}_{H,\gamma}$ is executed on such a dynamical system, the resulting system is $\gamma^{2}$-stable.
9.4. Prove Lemma 8.5.
9.5. Prove Lemma 8.7.
Chapter 10 Online Nonstochastic System Identification
A fundamental difficulty in applied control is that the dynamics governing the system behavior may be unknown, or too complex to describe. A prominent example is that of control in medical ventilators described in section 2.1.1. The respiratory system of a patient is significantly more complex than the fluid dynamics equations that are commonly used to describe it.
The subfield in control theory that deals with learning a given dynamical system from experience is called system identification. It is notoriously difficult to learn non-linear dynamical system well enough to be able to control them, much less so in high dimensions. This is further complicated in our treatment of control: discrete time representation, finite time performance guarantees, and adversarial perturbations as well as noise.
It is for this reason that we focus on the basic setting of linear time invariant dynamical systems. Thus, for the remainder of this chapter we consider dynamics of the form
$\mathbf{x}_{t+1}=A\mathbf{x}_{t}+B\mathbf{u}_{t}+\mathbf{w}_{t}.$
The question we ask in this chapter is: can we attain sublinear policy regret vs. the class of disturbance action policies, when the only information available to the learner are the sequential state observations $\mathbf{x}_{t}$? This may seem very challenging, since the perturbations are adversarially chosen.
Of course, for the problem to be feasible at all the system has to at least be stabilizable. We start with an even stronger assumption, suppose that the linear dynamical system is stable, is it learnable within the framework of online nonstochastic control? Perhaps surprisingly, the answer is positive in a strong sense.

### v4

@@ 7.5 Exercises
7.2.
  1. 1.
    Let $l:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a $\beta$-smooth function and let $A\in\mathbb{R}^{d\times n}$ be a matrix. Show that $l(A)$ is necessarily a smooth function for $x\in\mathbb{R}^{n}$.
  2. 2.
    Let $l:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a $\mu$-strongly convex function and let $A\in\mathbb{R}^{d\times n}$ be a matrix. Show that $l(A)$ is not necessarily strongly-convex for $x\in\mathbb{R}^{n}$.
7.3. Prove that the loss functions $\ell_{t}$ are convex in $M_{1:h}$, as per Lemma 7.3, even if the initial state is non-zero.
7.4. Consider control of a general, non-linear dynamical system. Give a version of the GPC algorithm 4 that can be applied to such a setting (without proving any performance guarantee).
7.5. In this exercise we prove and extend Theorem 7.2 for linear time-varying systems that are stabilizable, rather than stable.
  1. 1.
    Write down and prove the theorem for time-varying systems $\{A_{t},B_{t}\}$ that are all $\gamma$-stable.
  2. 2.
    Consider a sequence of $\gamma$-stabilizable time-varying systems $\{A_{t},B_{t}\}$. Show how to reduce this case to the previous by considering a sequence of linear time-varying $\gamma$-stable systems $\tilde{A}_{t},B_{t}$, where $\tilde{A}_{t}=A_{t}+B_{t}K_{t}$.
Chapter 8 Online Nonstochastic System Identification
In the previous chapters, we focused on control algorithms that assume knowledge of the system dynamics. However, in many real-world applications, system parameters are either unknown or change over time. This necessitates an approach where the controller must simultaneously learn the system model while making control decisions. This problem, known as system identification, is fundamental to adaptive control, reinforcement learning, and modern data-driven methods for control.
8.1 The Role of System Identification in Online Control
System identification plays a crucial role in control design, as it allows controllers to estimate unknown system parameters in real time. Unlike classical identification methods that rely on extensive offline data collection, an online approach must identify the system while actively controlling it. This presents several unique challenges:
  • Feedback loops and nonstationarity: Unlike supervised learning settings, where data is collected passively, control decisions influence future observations, making the estimation process highly dynamic.
  • Adversarial or arbitrary disturbances: The disturbances affecting the system may not follow a probabilistic model but can be chosen adversarially, leading to challenges in maintaining robustness.
  • Balancing exploration and performance: The controller must explore sufficiently to estimate system parameters accurately while ensuring stable performance, a tradeoff that directly impacts regret minimization.
8.2 A Nonstochastic Approach to System Identification
Traditional system identification techniques often rely on stochastic assumptions, where noise is assumed to be drawn from a known distribution, and system parameters are estimated using statistical inference techniques. However, in nonstochastic system identification, no assumptions are made about the noise distribution or system perturbations. Instead, identification strategies are designed to be robust to worst-case disturbances and operate under adversarial uncertainty.
This shift in perspective has several implications:
  • Performance guarantees without probabilistic assumptions: Unlike stochastic estimation methods, which provide guarantees under specific noise models, nonstochastic approaches aim for regret bounds that hold in arbitrary environments.
  • Regret as a measure of learning efficiency: The effectiveness of an identification strategy is measured not in terms of statistical efficiency but in its ability to achieve low regret, ensuring that the system is identified quickly enough to enable near-optimal control.
  • Computational tractability in identification: Classical system identification methods often involve solving nonconvex optimization problems, making them computationally demanding. Nonstochastic approaches frequently leverage convex relaxation techniques to sidestep intractability.
The problem of online system identification is closely related to online convex optimization and sequential decision-making. Many regret-minimization techniques developed for online learning can be adapted to system identification settings. A key distinction, however, is that in control settings, the learned system model directly affects future decisions, creating a coupling between estimation and control that does not typically exist in standard online learning formulations.
This connection between system identification and online learning leads to several fundamental questions:
  • How should a controller optimally trade off exploration (learning system parameters) and exploitation (minimizing control cost)?
  • What are the best algorithmic strategies for estimating system parameters while controlling in adversarial settings?
  • What regret guarantees can be achieved in online nonstochastic system identification, and how do they compare to traditional learning-based control approaches?